Using the following standard reduction potentials:

Ag+ (aq) + e- ⇌ Ag (s) Eo = +0.80 V

Pb2+ (aq) + 2e- ⇌ Pb (s) Eo = -0.13 V


i. Write a balanced redox reaction that represents the overall cell reaction for the feasible galvanic cell.


ii. Calculate the standard cell potential.


iii. Draw a fully labelled galvanic cell for this redox system. Your labels should show (but not limited to): the electron and ion flow; the anode and cathode, the half cells where oxidation and reduction take place.

1
Expert's answer
2021-06-22T04:42:00-0400

I) Pb(S) +2Ag+(aq)\to→ Pb2+(aq) +Ag(S)


(ii)Ag+ (aq)+ e-⇌ Ag (S) Eo= +0.80 V -Cathode


Pb2+ (aq) + 2e-⇌ Pb(S) Eo= -0.13 V -Anode


Eocell=Eocathode-Eoanode

=0.80V-(-0.13)V=0.93V


Standard cell potential =0.93V


(iii)


Electroplating is the process of applying a layer of metal or non-metal onto a metallic surface usually for decorative purposes or for protecting them from corrosion. The electrolysis process is employed for plating the metal.

In silver plating, the object to be plated i.e. the coin is made from the cathode of an electrolytic cell. While the anode is the silver bar and the electrolyte used is silver cyanide (AgCN). When current is passed through the cell, positively charged silver ions (Ag+) from the silver cyanide migrate to the negative anode (the coin). These ions are neutralized by electrons and stick to the coin as silver metal.

The anode bar dissolves to replenish silver ions in the solution which is transferred to the coin until a desired coating thickness is built-up on the coin surface.